Set up the integral that would give the area that is bounded between these functions. y = ĝ x = y2 – 4y A = ---Select--- A
Answer #1

PAGE NO - Hirst we find the Intexzeden beints of one - y=x and x=y= 44. i- c. x=34 and a= y²qy put 34 - 4² 44 = y ²44-34 =0 s